OpenStudy (anonymous):

(2x+1)^((1)/(2))=x-7 how would I solve for x?

OpenStudy (whpalmer4):

\[(2x+1)^{1/2} = x-7\] First thing to recognize is that \[u^{1/2} = \sqrt{u}\](and in general \(u^{m/n} = \sqrt[n]{u^m}\))

OpenStudy (whpalmer4):

So we could rewrite the problem as \[\sqrt{2x+1} = x-7\]and after the last several questions you've asked, I bet you know how to proceed with that, right? :-)

OpenStudy (anonymous):

Yes, thank you! :D